OpenStudy (anonymous):

points B,D and F are midpoints of the sides of ACE EC=32 and BF=2x-10 find the value of x

OpenStudy (anonymous):

BF will be half of EC so BF = 1/2* 32 = 16 so 2x - 10 = 16 2x = 16 + 10 x = 26/2 = 13

OpenStudy (anonymous):

This based on the fact that a line joining midpoints of two sides or any triangle is parallel to the third side and half of it.....